# CCAT: Mod-Cam Readout Overview and Flexible Stripline Performance

### v1

@@ IV Crosstalk
Figure 6: Left: The connectorized end of one stripline circuit. The numbers denote the labeling scheme used in Fig. 7 top and middle panels. Right: Photograph of the SMP-SMA transition PCB. Two PCBs plug into each end of the stripline. The numbers denote the labeling scheme used in the bottom panel of Fig. 7.
@@ IV-A Measurement
During in-lab testing of the 280 GHz instrument module we observed elevated crosstalk through the full readout chain. Cold measurements confirmed no significant reflections in the readout chains, so we are confident excess pickup is primarily occurring between different readout chains. We sometimes refer to this as inter-channel crosstalk. In order to isolate the source of this crosstalk, we perform warm S21 measurements using an Agilent E5071C VNA over the full readout bandwidth for the TiN and Al arrays installed in the 280 GHz module, which combined span a range from 300 MHz - 1.5 GHz. Crosstalk measurements are performed for a variety of configurations: a lone stripline, a stripline with a SMP-SMA transition PCB connected on either end (mimicking the readout structure in Mod-Cam), and a lone transition PCB.
@@ IV-B Crosstalk Measurement Results
Figure 7: Warm crosstalk measurements are shown. The top panel shows the crosstalk measured for all channel combinations with a single channel (6) in an isolated stripline. The labeling scheme indicates the Input (I) and Output (O) with each associated channel number. Transmission through channel 6 is observed at the top of the plot. The middle panel shows crosstalk measured through the combination of a stripline with a SMP-SMA transition PCB on either end. The bottom panel shows crosstalk measured through an isolated transition PCB.
@@
Warm S21 measurements of the various configurations are shown in Fig. 7, with labels corresponding to connections designated in the labeling scheme in Fig. 6. The results demonstrate that the stripline alone is not the dominant source of crosstalk in the readout chain, with inter-channel crosstalk 50 dB or more below transmission, meeting project requirements. The middle panel of Fig. 7 isolates the source of crosstalk in the readout chain to the custom transition PCBs; in this configuration, chains 2, 4, and 6 share the same PCB and show elevated crosstalk, up to 30 dB above those combinations that do not share a PCB. The level of crosstalk is also correlated with physical spacing on the transition PCB, with closer traces observing higher pickup; this accounts for the elevated level of the $\mathrm{6I\_4O}$” measurement relative to the $\mathrm{6I\_2O}$” measurement.
In light of this result, the SMP-SMA transition PCBs have been redesigned using an impedance matched via to quickly bury the signal trace in an intermediate PCB layer between ground planes for improved isolation. Given the frequencies of the readout band, radiative effects from the via ends are expected to be minimal. Moreover, our ability to isolate signal traces on the PCB is fundamentally limited by the ganged SMP connector design, which is necessarily mounted on the edge of the PCB board. At minimum, this results in a small section of microstrip trace on the PCB which can add crosstalk. The original PCB design referenced in Fig. 6 and Fig. 7 utilized only microstrip traces.
@@
To estimate the predicted crosstalk of this new design we performed simulations of the microstrip-to-stripline transition via structure using Ansys HFSS software. Using these simulations we first optimize the microstrip, stripline, and via dimensions to impedance match the full structure to $50~\Omega$. Using a simulated PCB configuration with optimal trace and via dimensions, the simulations yield crosstalk estimates at a level no greater than -60 dB for a single structure, and -57 dB across the full PCB. This represents an approximately 20 dB improvement in crosstalk level over the old PCB design, which should allow the PCB to exceed project requirements.
@@ V Conclusion
We performed two types of measurements to characterize the thermal loading through the stripline flexible circuitry designed for Mod-Cam and Prime-Cam: finding the $RRR$ of the copper traces and directly estimating the thermal conductivity of the circuit as a function of temperature. The measured thermal conductivity values are consistent with theoretical values common in literature. With the materials of the stripline characterized, we then predicted the expected thermal loading for nominal stripline configurations for both Mod-Cam and Prime-Cam. Results of these predictions demonstrate the significant improvement in expected loading levels for the redesigned Prime-Cam stripline; despite being populated with three times more stripline, the estimated loading at 4 K through one Prime-Cam readout harness is actually lower than the Mod-Cam readout harness.
When in-lab testing of the readout in Mod-Cam yielded higher-than-anticipated crosstalk through the readout chains, we performed measurements to isolate the source of this pickup. Results of this testing identified the source to be the SMP-SMA transition PCBs, which carry signals from the stripline flexible circuits into coax cables. Subsequent redesign and simulation of these PCBs has demonstrated that they should meet project requirements for low crosstalk when fabricated.
Flexible stripline circuits installed in Mod-Cam have proven to be a robust option for the simultaneous readout of thousands of KIDs during months of in-lab testing. The results presented in this paper validate stripline as a scalable solution for readout of large KID arrays and inform the ongoing fabrication of optimized Prime-Cam flexible circuit assemblies.

### v2

@@ IV-A Measurement
During in-lab testing of the 280 GHz instrument module we observed elevated crosstalk through full readout chains. Cold measurements confirmed no significant reflections in the readout chains, so we are confident excess pickup is primarily occurring between different readout chains. We sometimes refer to this as inter-channel crosstalk. In order to isolate the source of this crosstalk, we performed warm S21 measurements using an Agilent E5071C VNA over the full readout bandwidth for the TiN and Al arrays installed in the 280 GHz module, which combined span a range from 300 MHz - 1.5 GHz. Crosstalk measurements were performed for a variety of configurations: a lone stripline, a stripline with a SMP-SMA transition PCB connected on either end (mimicking the readout structure in Mod-Cam shown in Fig. 2), and a lone transition PCB.
@@ IV-B Crosstalk Measurement Results
Figure 6: Warm crosstalk measurements are shown. Top: Crosstalk measured for all channel combinations with a single channel (6) on an isolated stripline. The labeling scheme, following that shown in Fig. 2, indicates the Input (I) and Output (O) with each associated channel number. All inter-channel measurements show comparable S21 amplitude, while transmission through channel 6 is observed at the top of the panel. Middle: Crosstalk measured through the combination of a stripline with SMP-SMA transition PCBs on either end. The two elevated traces with uniform standing wave patterns believed to result from connectors correspond to traces that share a connector, and hence also an SMP-SMA transition PCB. Bottom: Crosstalk measured through an isolated transition PCB following the PCB labeling scheme in Fig. 2. Channels spaced physically further on the PCB show lower S21 amplitude than those immediately adjacent.
@@
Warm S21 measurements of the configurations are shown in Fig. 6, with labels corresponding to connections designated in the labeling scheme in Fig. 2. The results demonstrate that the stripline alone is not the dominant source of crosstalk in the readout chain, with inter-channel crosstalk shown in the top panel of Fig. 6 uniformly 50 dB or more below transmission, meeting project requirements. Additional structure visible in the inter-channel traces on this plot are primarily attributed to the SMP connections between the stripline and PCBs; this structure was observed to shift slightly when connectors were manually perturbed.
The middle “stripline + PCB” panel of Fig. 6 isolates the source of crosstalk in the readout chain to the custom transition PCBs; in this configuration, chains 2, 4, and 6 share the same PCB and show elevated crosstalk, up to 30 dB above those combinations that do not share a PCB. The level of crosstalk is also correlated with physical spacing on the transition PCB, accounting for the elevated level of the $\mathrm{6I\_4O}$ measurement relative to the $\mathrm{6I\_2O}$ measurement.
Inter-channel crosstalk measurements in the “PCB-only” panel of Fig. 6 also confirm the SMP-SMA transition PCB as the source of elevated crosstalk. The crosstalk in trace $PCB\_\mathrm{3I\_2O}$ for a single PCB is 40dB-40~dB at 1.4 GHz. For the full chain two PCBs are used, resulting in the 3dB3~dB higher baseline crosstalk demonstrated in the “stripline + PCB” panel of Fig. 6 in the corresponding trace labeled $\mathrm{6I\_2O}$. In light of this result, the SMP-SMA transition PCBs have been redesigned using an impedance matched via to quickly bury the signal trace in an intermediate PCB layer between ground planes for improved isolation. Radiative effects from the via ends are expected to be minimal from signals in the readout bandwidth.
@@
To estimate the predicted crosstalk of the new PCB design we performed simulations of the via-trace structure on the PCB using Ansys HFSS software, with critical dimensions selected to impedance match the full structure to $50~\Omega$. These simulations estimate crosstalk at a level no greater than -60 dB for a single structure, and -57 dB across the full PCB. This represents an approximately 20 dB improvement in crosstalk level over the prior PCB design, allowing the PCB to exceed project requirements.
@@ V Conclusion
We performed two types of measurements to characterize the thermal loading through the stripline flexible circuitry designed for Mod-Cam and Prime-Cam: finding the $RRR$ of the copper traces and directly estimating the thermal conductivity of the circuit as a function of temperature for various temperature ranges. Using the $RRR$ result, we predicted the expected thermal loading for nominal stripline configurations for both Mod-Cam and Prime-Cam. Results of these predictions demonstrate the significant improvement in expected loading for the redesigned Prime-Cam stripline circuits.
When in-lab testing of the readout in Mod-Cam yielded higher-than-anticipated crosstalk through the readout chains, we performed measurements to isolate the source of this pickup. Results of this testing identified the source as the SMP-SMA transition PCBs, which carry signals from the stripline flexible circuits into coaxial cables. Subsequent redesign and simulation of these PCBs has demonstrated that they should meet project requirements for low crosstalk when fabricated. This new PCB design should be implemented in both Mod-Cam and Prime-Cam.
Flexible stripline circuits installed in Mod-Cam have proven to be a robust option for the simultaneous readout of thousands of KIDs during months of in-lab testing. The results presented in this paper validate stripline as a scalable solution for readout of large-format KID arrays and inform the ongoing fabrication of optimized Prime-Cam flexible circuit assemblies.